squareroot(2) times cuberoot(3) in radical form
Answer by stanbon(75887)

squareroot(2) times cuberoot(3) in radical form
-----
2^(1/2)* 3^(1/3)
----
= 2^(3/6)*3^(2/6)
----
= (2^3*3^2)^(1/6)
----
= (8*9)^(1/6)
------
= 6th root of 72
======================
Cheers,
Stan H.